OpenStudy (anonymous):

Identify the degree of the following polynomial: 8x – x3 – 3x – 12x9 – 6x2

OpenStudy (anonymous):

8x – x3 – 3x – 12x9 – 6x2 = -12x^9 -x^3 -6x^2 +8x -3x = -12x^9 -x^3 -6x^2 +5x Highest power is 9 Degree = 9